OpenStudy (quinton1011):

Which terms are like terms? 7q, 9q5, 2q, 28p3, 13p2, 25q5, 19p2, and 32p3 Choose all answers that are correct. A. 7q and 2q B. 13p2 and 19p2 C. 28p3 and 32p3 D. 9q5 and 25q5

OpenStudy (anonymous):

C. 28p3 and 32p3

OpenStudy (quinton1011):

there can be more than one answer, is there?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

D. 9q5 and 25q5 too

OpenStudy (anonymous):

no others

OpenStudy (quinton1011):

okay thanks here's a medal

OpenStudy (quinton1011):

could you help me with a few more?

OpenStudy (quinton1011):

?

OpenStudy (anonymous):

sure

OpenStudy (quinton1011):

okay thanks

OpenStudy (quinton1011):

Simplify. −1w + 4w4 + y4 – 3y4 – 2w – 3w4 A. −2w4 – 2y4 B. −w4y4 – 3w C. w4 – 3w – 2y4 D. 12w4 + 2w – 3y4

OpenStudy (anonymous):

is it multiply choice

OpenStudy (quinton1011):

yeah, there's only one correct one

OpenStudy (anonymous):

k i think i got the answer one sec let me check my self

OpenStudy (quinton1011):

okay :)

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

next

OpenStudy (quinton1011):

Bill and Ron build towers using blocks. The expression 4b + 2 represents the number of blocks Bill uses. The expression 8b – 5 represents the number of blocks Ron uses. Complete the expression to represent the total number of blocks Bill and Ron use. b −

OpenStudy (anonymous):

brb

OpenStudy (quinton1011):

i'v only got 5 min

OpenStudy (quinton1011):

4 min

OpenStudy (milanitaly1618):

So you have to add both of the expressions that both bill and Ron have so it would be....

OpenStudy (milanitaly1618):

4b + 2 + 8b – 5 and 4b and 8b are like terms do you'd add them

OpenStudy (quinton1011):

and milan can you give epicgammer a medal

OpenStudy (milanitaly1618):

ok

OpenStudy (quinton1011):

i gave you a medal

OpenStudy (milanitaly1618):

Thanks

OpenStudy (quinton1011):

i'm still a little confused on this

OpenStudy (milanitaly1618):

So you have to combine the like terms right so since 4b + 2 and 8b – 5 both have 4b and 8b you combine them to get 12b-7

OpenStudy (quinton1011):

oh okay thank you

OpenStudy (anonymous):

im back

OpenStudy (milanitaly1618):

No problem

OpenStudy (quinton1011):

i'v got two more questions

OpenStudy (milanitaly1618):

ok

OpenStudy (anonymous):

k ask them

OpenStudy (quinton1011):

A teacher wrote expressions on cards and put them in a bag. Thomas and Kelly drew cards from the bag and added the expressions. Thomas drew the expressions 3x + 7 and 2x – 3. Kelly drew the expressions −6x + 3 and 14x + 2. How much greater was Kelly’s sum than Thomas’s sum? x +

OpenStudy (anonymous):

let me think

OpenStudy (milanitaly1618):

Well first you would have to combine like terms for both Kelly and Thomas.

OpenStudy (quinton1011):

okay

OpenStudy (milanitaly1618):

Thomas=3x + 7+ 2x – 3= 5x-10 Kelly=−6x + 3+ 14x + 2= 8x+5

OpenStudy (quinton1011):

okay...

OpenStudy (milanitaly1618):

Then is since you dont know what x equals you can put an number there to estimate which is greater 5x-10 8x+5 we can make x=9 45-10=35 thomas 72+5=77 kelly

OpenStudy (milanitaly1618):

Sorry it took me so long i'm also in a session

OpenStudy (quinton1011):

i was late for science :(

OpenStudy (milanitaly1618):

Yeah I saw did that help your question?

OpenStudy (quinton1011):

yeah i have one more

OpenStudy (milanitaly1618):

ok

OpenStudy (quinton1011):

What is the perimeter of a triangle with the given side lengths? s1 = x – 3 ft s2 = x + 2 ft s3 = x + 4 ft P = x + ft

OpenStudy (quinton1011):

i suck at math

OpenStudy (milanitaly1618):

I'm sorry but I'm not exactly sure how to do this one:(

OpenStudy (quinton1011):

:'( okay thanks for helping

OpenStudy (milanitaly1618):

No problem I'm glad I was able to help with your other questions:)

OpenStudy (quinton1011):

i'll get someone else to help with this one :)

OpenStudy (milanitaly1618):

Yeah

OpenStudy (quinton1011):

yeah

OpenStudy (quinton1011):

my computer crashed

OpenStudy (milanitaly1618):

not good

OpenStudy (quinton1011):

I'm gonna fail this......

OpenStudy (milanitaly1618):

No you wont you can post it as a new question:)

OpenStudy (quinton1011):

i'm closing this one

OpenStudy (milanitaly1618):

ok

OpenStudy (anonymous):

yeah here take 2+4+3 and then take the 9 and thats x plus 9

OpenStudy (anonymous):

you can close it and start up a new one if you want